A circular loop of wire of radius 0.25 cm is expected to an external magnetic field that is perpendicular to the plane of the loop. For a time, the magnetic field decreases at a rate of 0.0035 T/s. What is the magnitude of the induced electromagnetic force in the loop?


Correct the given Solution:


Solution:

F= SB cos α,

S =πr²,  α= 0°

F= SB cos α

 or F=πr² B

F = (3.14) (0.25) ²  (0.0035)

F= 0.000686875

Or

F= 6.9 x 10-8 N

1
Expert's answer
2022-05-25T14:25:23-0400
"{\cal E}=-\frac{\Delta \Phi}{\Delta t}=-S\frac{\Delta B}{\Delta t}=-\pi r^2\frac{\Delta B}{\Delta t}"

"{\cal E}=-3.14*(0.25*10^{-2})^2*(-0.0035)\\
=6.9*10^{-8}\:\rm V"